# Password Strength Detection via Machine Learning: Analysis, Modeling, and Evaluation

### v1

@@ 1. INTRODUCTION
The insecure practices of users selecting passwords have significantly compromised their effectiveness, leading to passwords that often fail to meet theoretical strength standards (Adams and Sasse, 1999). Common issues include the use of overly simplistic passwords and the tendency to reuse passwords across multiple accounts (Dell’Amico et al., 2010). This practice of password reuse exposes user accounts to ”crash attacks,” (Singer et al., 2013) thereby providing attackers with greater opportunities to exploit vulnerabilities. Moreover, many web application service providers continue implementing outdated password security measures that are ill-equipped to combat evolving attack, resulting in significant data leaks and exacerbating the password security crisis (Florêncio and Herley, 2010). Consequently, it is imperative to investigate strategies for enhancing password security and mitigating various network attacks to protect personal privacy and uphold the security of network systems.
@@ 2.1. Password Cracking Attack
Hybrid Attacks. Hybrid attacks employ a combination of multiple password-cracking methodologies previously delineated (Narayanan and Shmatikov, 2005). For example, an attacker may initially attempt to penetrate common passwords utilizing a dictionary attack, subsequently employing brute force on remaining passwords that remain unbroken. This approach capitalizes on the strengths of diverse password-cracking methods, thereby enhancing the overall likelihood of success. Hybrid attacks may also encompass rule-based tactics, whereby the attacker generates password combinations predicated on specific rules, such as appending numbers or special characters to a dictionary list. This strategy effectively addresses the countermeasures users may implement to fortify their passwords, thus increasing the probability of a successful breach.
@@ 2.2. Password Defense
Periodic Password Changes. Implementing periodic password changes represents an effective defense strategy aimed at mitigating risks associated with prolonged use of the same password. When users are encouraged to update their passwords regularly, the potential window of opportunity for attackers to access protected resources is minimized, even in instances where an old password is compromised. Organizations should establish clearly defined password change intervals that mandate employees to refresh their passwords at specific times (Bellovin and Merritt, 1992). However, these intervals should not be excessively brief, as this may lead to the selection of simplistic passwords or instances of forgotten credentials.
@@
Password Hashing and Salting. In the context of password storage, employing encryption through a hashing function is a widely accepted security practice. This function transforms the password into a fixed-length hash value, rendering it infeasible for an attacker to recover the original plaintext password (Shay et al., 2010). To further bolster security, the practice of salting can be employed, wherein random characters are appended to both the beginning and end of the password prior to hashing. This salting technique effectively mitigates the risk of rainbow table attacks and precomputation attacks that target the same hashing function.
@@ 2.3. Application of Machine Learning to Password Security
Password Cracking Predictions. Through the utilization of machine learning, security experts can anticipate the strategies and passwords that attackers are likely to target. By training and analyzing models based on a substantial dataset of cracking incidents, researchers can expose the tactics and behavioral traits exhibited by attackers. Machine learning algorithms facilitate the extraction of all pertinent parameters required for classifying various datasets associated with attacks. This intelligence can be leveraged to devise targeted defense strategies, significantly diminishing the likelihood of successful attacks.
Defense Strategy Optimization. Machine learning can optimize existing password defense strategies, thereby enhancing security and reliability. The application of machine learning algorithms within the field of cryptography is extensive (Zou et al., 2025). These algorithms can serve as tools for assessing the effectiveness of multi-factor authentication methods, thus assisting organizations in selecting the authentication protocols that best align with their specific requirements. Furthermore, machine learning can identify and remediate vulnerabilities within security configurations, such as instances involving devices operating with default passwords, ultimately improving the overall security posture.
@@
Intelligent Password Generator. Machine learning can be utilized to create intelligent password generators that produce robust and memorable passwords tailored to user requirements and security specifications (Li et al., 2014). By training a model to comprehend the structure of human language and lexical patterns, such a generator is capable of crafting passwords that satisfy specific strength criteria while preserving an element of readability. This methodology enhances password security and mitigates the likelihood of users forgetting their passwords.
@@ 3. DATA COLLECTION AND PREPROCESSING
This research employs real user data publicly leaked from networks to analyze the password database. We collected publicly leaked user information datasets from over 20 websites, including notable platforms such as JD, Taobao, and CSDN, yielding a total user data volume exceeding one billion records. These datasets have been compromised through various means, including internal leaks by website administrators and external breaches perpetrated by malicious actors who exploit vulnerabilities in social software, email systems, and institutional infrastructures. Such breaches often involve using social engineering or weak password protocols to secure unauthorized database access. This section selects 11 representative user datasets based on website types, categorized into lifestyle shopping, social forums, and video entertainment, with detailed sources in Table 1.
@@ 4 Mop Forum 8731286 5337012 E-mail, Password
6 7k7k Game 19119973 4930689 E-mail, Password
@@ 4. DATA CHARACTERIZATION
This section will thoroughly analyze the characteristics of password data sources, including analysis of familiar patterns, password length, and character composition, to uncover users’ password selection habits and vulnerabilities across different websites. By comparing the strength and characteristics of passwords from each data source, this study aims to provide tailored password security recommendations and protective measures for enterprises and individuals.
@@ 4.1. Familiar Patterns
A ranking of the top ten most prevalent passwords has been compiled by sorting them based on their frequency of occurrence in the user datasets, with results presented in Table 3. The proportion of these top ten passwords ranges from 3.30% (Shopping 1) to 10.76% (Forum 1). This variance suggests that distinct website types and user demographics exhibit differing preferences in password selection. Based on the analyzed password patterns, categories have been delineated as follows: sequential numbers, repeated characters, keyboard patterns, simple fragment combinations, and semantically significant passwords.
@@
    Special Semantic Passwords: Passwords like ”5201314” and ”1314520”, which carry special semantic meaning, suggest that users might select personally significant terms. Examples such as ”qiulaobai”, ”111222tianya”, and ”dearbook” may connect to specific websites or cultural references. While these passwords could be more challenging to guess, they remain vulnerable to targeted attacks if an assailant conducts a thorough analysis of the relevant cultural context or website content.
@@
2 111111 111111 12345678 a123456 123456789 123456789
3 123456789 zz12369 $11$ 123456789 111111 111111
4 123123 qiulaobai dearbook 111111 123123 123123
5 111222tianya 123456aa $00$ 5201314 5201314 5201314
6 5201314 wmsxie123 123123123 123123 789456 12345
7 123321 123123 1234567890 a321654 666666 12345678
8 12345678 000000 $88$ 12345 123321 123321
9 666666 qq66666 111111111 000000 1314520 1314520
@@ 4.5. Common Attack Patterns and Threats
    Targeted attacks: Attackers may engage in focused efforts to guess passwords based on factors such as the distinct characteristics of the target site and user demographics. For instance, attempts may be made to guess semantic passwords that are culturally relevant to specific websites. These types of passwords are prevalent in the Forum 2 and Forum 3 datasets, indicating that targeted attacks may represent a considerable threat to these datasets.
@@ 5.2.1. Model Selection
Neural Networks: Neural networks, recognized for their capability as robust nonlinear classifiers, can effectively extract complex features (Zhong et al., 2023). Simple neural network architectures, such as multilayer perceptrons (MLP), are recommended for strong and weak password classification. Although requiring extended training durations, they typically produce superior classification outcomes.
@@ 5.3. Model Evaluation
Model evaluation represents the final phase following the training and validation of a model, with the objective of assessing the model’s performance on previously unseen data. In this stage, we evaluate the selected model utilizing a test set and compute various evaluation metrics, including accuracy, recall, and F1 score, to establish a benchmark for the model’s anticipated performance in real-world applications.
We evaluate model performance based on the test set for the six models, as detailed in Table 7.
@@ 6. DEFENCE STRATEGIES RECOMMENDATION
It is imperative to implement a variety of effective password policies and management strategies to enhance password security. The following recommendations are put forth:

### v2

@@ 1. INTRODUCTION
The insecure practices of users selecting passwords have significantly compromised their effectiveness, leading to passwords that often fail to meet theoretical strength standards (Adams and Sasse, 1999). Common issues include the use of overly simplistic passwords and the tendency to reuse passwords across multiple accounts (Dell’Amico et al., 2010) (Das et al., 2014). This practice of password reuse exposes user accounts to ”crash attacks,” (Singer et al., 2013) thereby providing attackers with greater opportunities to exploit vulnerabilities. Moreover, many web application service providers continue implementing outdated password security measures that are ill-equipped to combat evolving attack, resulting in significant data leaks and exacerbating the password security crisis (Florêncio and Herley, 2010). Consequently, it is imperative to investigate strategies for enhancing password security and mitigating various network attacks to protect personal privacy and uphold the security of network systems.
@@ 2.1. Password Cracking Attack
Hybrid Attacks. Hybrid attacks employ a combination of multiple password-cracking methodologies previously delineated (Narayanan and Shmatikov, 2005) (Shay et al., 2016). For example, an attacker may initially attempt to penetrate common passwords utilizing a dictionary attack, subsequently employing brute force on remaining passwords that remain unbroken. This approach capitalizes on the strengths of diverse password-cracking methods, thereby enhancing the overall likelihood of success. Hybrid attacks may also encompass rule-based tactics, whereby the attacker generates password combinations predicated on specific rules, such as appending numbers or special characters to a dictionary list. This strategy effectively addresses the countermeasures users may implement to fortify their passwords, thus increasing the probability of a successful breach.
@@ 2.2. Password Defense
Periodic Password Changes. Implementing periodic password changes represents an effective defense strategy aimed at mitigating risks associated with prolonged use of the same password. When users are encouraged to update their passwords regularly, the potential window of opportunity for attackers to access protected resources is minimized, even in instances where an old password is compromised (Hao and van Oorschot, 2022). Organizations should establish clearly defined password change intervals that mandate employees to refresh their passwords at specific times (Bellovin and Merritt, 1992). However, these intervals should not be excessively brief, as this may lead to the selection of simplistic passwords or instances of forgotten credentials.
@@
Password Hashing and Salting. In the context of password storage, employing encryption through a hashing function is a widely accepted security practice. This function transforms the password into a fixed-length hash value, rendering it infeasible for an attacker to recover the original plaintext password (Shay et al., 2010). To further bolster security, the practice of salting can be employed, wherein random characters are appended to both the beginning and end of the password prior to hashing. This salting technique effectively mitigates the risk of rainbow table attacks and precomputation attacks that target the same hashing function (Sriramya and Karthika, 2015).
@@ 2.3. Application of Machine Learning to Password Security
Password Cracking Predictions. Through the utilization of machine learning, security experts can anticipate the strategies and passwords that attackers are likely to target. By training and analyzing models based on a substantial dataset of cracking incidents, researchers can expose the tactics and behavioral traits exhibited by attackers (Weir et al., 2009). Machine learning algorithms facilitate the extraction of all pertinent parameters required for classifying various datasets associated with attacks. This intelligence can be leveraged to devise targeted defense strategies, significantly diminishing the likelihood of successful attacks.
Defense Strategy Optimization. Machine learning can optimize existing password defense strategies, thereby enhancing security and reliability. The application of machine learning algorithms within the field of cryptography is extensive (Zou et al., 2025) (Kelley et al., 2012). These algorithms can serve as tools for assessing the effectiveness of multi-factor authentication methods, thus assisting organizations in selecting the authentication protocols that align with their requirements. Furthermore, machine learning can identify and remediate vulnerabilities within security configurations, such as devices operating with default passwords, ultimately improving the overall security posture.
@@
Intelligent Password Generator. Machine learning can be utilized to create intelligent password generators that produce robust and memorable passwords tailored to user requirements and security specifications (Li et al., 2014b). By training a model to comprehend the structure of human language and lexical patterns, such a generator is capable of crafting passwords that satisfy specific strength criteria while preserving an element of readability. This methodology enhances password security and mitigates the likelihood of users forgetting their passwords.
@@ 3. DATA COLLECTION AND PREPROCESSING
Our study employs real user data publicly leaked from networks to analyze the password database. We collected publicly leaked user information datasets from over 20 websites, including notable platforms such as JD, Taobao, and CSDN, yielding a total user data volume exceeding one billion records (Bonneau, 2012) (Li et al., 2014a). These datasets have been compromised through various methods, including internal leaks by website administrators and external breaches perpetrated by malicious actors who exploit vulnerabilities in social software, email systems, and institutional infrastructures. Such breaches involve using social engineering or weak password protocols to secure unauthorized database access. This section selects 11 representative user datasets based on website types, categorized into lifestyle shopping, social forums, and video entertainment, with detailed sources in Table 1.
@@ 4 Mop Forum 8731286 5337012 E-mail, Password
6 $7k$ Game 19119973 4930689 E-mail, Password
@@ 4. DATA CHARACTERIZATION
This section will thoroughly analyze the characteristics of password data sources, including analysis of familiar patterns, password length, and character composition, to uncover users’ password selection habits and vulnerabilities across different websites (Florencio and Herley, 2007). By comparing the strength and characteristics of passwords from each data source, this study aims to provide tailored password security recommendations and protective measures for enterprises and individuals.
@@ 4.1. Familiar Patterns
A ranking of the top ten most prevalent passwords has been compiled by sorting them based on their frequency of occurrence in the user datasets, with results presented in Table 3. The proportion of these top ten passwords ranges from 3.30% (Shopping 1) to 10.76% (Forum 1). This variance suggests that distinct website types and user demographics exhibit differing preferences in password selection (Mittal and Ilavarasan, 2019). Based on the analyzed password patterns, categories have been delineated as follows: sequential numbers, repeated characters, keyboard patterns, simple fragment combinations, and semantically significant passwords.
@@
    Special Semantic Passwords: Passwords like ”5201314” and ”1314520”, which carry special semantic meaning, suggest that users might select personally significant terms. Examples such as ”qiulaobai”, ”111222tianya”, and ”dearbook” may connect to specific websites or cultural references. While these passwords could be more challenging to guess, they remain vulnerable to targeted attacks if an assailant conducts a thorough analysis of the relevant cultural context or website content (Mazurek et al., 2013).
@@
2 $111$ $111$ 12345678 a123456 123456789 123456789
3 123456789 zz12369 $1111$ 123456789 $111$ $111$
4 $123$ qiulaobai dearbook $111$ $123$ $123$
5 111222tianya 123456aa $0000$ 5201314 5201314 5201314
6 5201314 wmsxie123 123123123 $123$ 789456 12345
7 123321 $123$ 1234567890 a321654 $666$ 12345678
8 12345678 $000$ $8888$ 12345 123321 123321
9 $666$ qq66666 111111111 $000$ 1314520 1314520
@@ 4.5. Common Attack Patterns and Threats
    Targeted attacks: Attackers may engage in focused efforts to guess passwords based on factors such as the distinct characteristics of the target site and user demographics (Hitaj et al., 2019). For instance, attempts may be made to guess semantic passwords that are culturally relevant to specific websites. These types of passwords are prevalent in the Forum 2 and Forum 3 datasets, indicating that targeted attacks may represent a considerable threat to these datasets.
@@ 5.2.1. Model Selection
Neural Networks: Neural networks, recognized for their capability as robust nonlinear classifiers, can effectively extract complex features (Zhong et al., 2023) (Kim et al., 2023). Simple neural network architectures, such as multilayer perceptrons (MLP), are recommended for strong and weak password classification. Although requiring extended training durations, they produce superior classification outcomes.
@@ 5.3. Model Evaluation
Model evaluation represents the final phase following the training and validation of a model, with the objective of assessing the model’s performance on previously unseen data (De Carnavalet et al., 2014). In this stage, we evaluate the selected model utilizing a test set and compute various evaluation metrics, including accuracy, recall, and F1 score, to establish a benchmark for the model’s anticipated performance in real-world applications. We evaluate model performance based on the test set for the six models, as detailed in Table 7.
@@ 6. DEFENCE STRATEGIES RECOMMENDATION
It is imperative to implement a variety of effective password policies and management strategies to enhance password security (Chiasson et al., 2008). The following recommendations are put forth: